# H-RDT: Human Manipulation Enhanced Bimanual Robotic Manipulation

### v1

@@ Introduction
  • Scratch: Our model without human pre-training (training from scratch)
@@
Performance Level RDT Scratch H-RDT
@@
Performance Level RDT Scratch H-RDT
@@
For the towel folding task (Table 1), H-RDT achieves a 52% complete success rate compared to 40% for RDT and 0% for training from scratch. The scratch model fails to achieve any complete folding, showing only partial success at lower skill levels, while RDT and H-RDT demonstrate more sophisticated manipulation capabilities.
The cup-to-coaster task (Table 2) shows H-RDT achieving a 64% complete success rate compared to 28% for RDT and 20% for scratch training. H-RDT demonstrates the lowest failure rate and fewer instances of partial success, indicating more robust performance for tasks requiring spatial reasoning to select appropriate arms.
Overall, H-RDT achieves an average success rate of 58% across both bimanual tasks compared to 34% for RDT and 10% for scratch training, demonstrating the effectiveness of human manipulation priors for diverse coordination challenges including deformable object manipulation and spatial reasoning tasks.
@@
The EgoDex pretraining dataset contains extensive pick-and-place manipulation patterns similar to these tasks, providing rich prior knowledge about how to perform such operations. Under these challenging conditions—where even state-of-the-art models like $\boldsymbol{\pi}_{0}$ struggle to properly fit the limited demonstration trajectories—H-RDT’s human manipulation priors enable noticeable performance improvements. H-RDT achieves an average success rate of 41.6% compared to 16.0% for RDT, 31.2% for $\boldsymbol{\pi}_{0}$, and 17.6% for Scratch, demonstrating the value of human manipulation priors for few-shot learning in data-limited scenarios.
Task Category RDT $\boldsymbol{\pi}_{0}$ Scratch H-RDT
@@
Subtask RDT $\boldsymbol{\pi}_{0}$ Scratch H-RDT
@@
H-RDT achieves the highest average success rate of 68.7% in Easy mode and 25.6% in Hard mode, demonstrating significant improvements over existing methods. H-RDT substantially surpasses training from scratch (Scratch) by 8.4% in both Easy and Hard modes, validating the effectiveness of human manipulation pre-training.
@@
Task RDT $\boldsymbol{\pi}_{0}$ Scratch H-RDT
@@
In multi-task settings, H-RDT achieves an average success rate of 87.2%, significantly outperforming RDT (28.8%), $\boldsymbol{\pi}_{0}$ (48.4%), and Scratch (67.2%). H-RDT shows a substantial 20.0% absolute improvement over training from scratch, which is notably larger than the improvements observed in single-task scenarios, demonstrating that human manipulation pre-training provides even greater advantages when learning across diverse tasks simultaneously.
@@
Impact of Human Pre-training: The consistent improvements of H-RDT over Scratch(H-RDT w/o human) across all experimental settings validate our core hypothesis that human manipulation data provides valuable inductive biases. The benefits are most pronounced in few-shot real-world scenarios, where human priors about object affordances and manipulation strategies prove crucial.
@@ Conclusion
Comprehensive evaluations across both simulation and real-world scenarios demonstrate consistent improvements over state-of-the-art methods, validating that human manipulation priors provide powerful inductive biases for sample-efficient robotic learning.
@@ A.2 Franka-Panda Detailed Results
Task RDT $\boldsymbol{\pi}_{0}$ Scratch H-RDT
@@ B.4 Real-World Task Definitions
Task ACT DP RDT $\boldsymbol{\pi}_{0}$ Scratch H-RDT

### v2

@@ Introduction
  • w/o human: Our model without human pre-training
@@
Performance Level RDT w/o human H-RDT
@@
Performance Level RDT w/o human H-RDT
@@
For the towel folding task (Table 1), H-RDT achieves a 52% complete success rate compared to 40% for RDT and 0% for training from scratch. The model without human data fails to achieve any complete folding, showing only partial success at lower skill levels, while RDT and H-RDT demonstrate more sophisticated manipulation capabilities.
The cup-to-coaster task (Table 2) shows H-RDT achieving a 64% complete success rate compared to 28% for RDT and 20% for training without human data. H-RDT demonstrates the lowest failure rate and fewer instances of partial success, indicating more robust performance for tasks requiring spatial reasoning to select appropriate arms.
Overall, H-RDT achieves an average success rate of 58% across both bimanual tasks compared to 34% for RDT and 10% for training without human data, demonstrating the effectiveness of human manipulation priors for diverse coordination challenges including deformable object manipulation and spatial reasoning tasks.
@@
The EgoDex pretraining dataset contains extensive pick-and-place manipulation patterns similar to these tasks, providing rich prior knowledge about how to perform such operations. Under these challenging conditions—where even state-of-the-art models like $\boldsymbol{\pi}_{0}$ struggle to properly fit the limited demonstration trajectories—H-RDT’s human manipulation priors enable noticeable performance improvements. H-RDT achieves an average success rate of 41.6% compared to 16.0% for RDT, 31.2% for $\boldsymbol{\pi}_{0}$, and 17.6% for H-RDT w/o human, demonstrating the value of human manipulation priors for few-shot learning in data-limited scenarios.
Task Category RDT $\boldsymbol{\pi}_{0}$ w/o human H-RDT
@@
Subtask RDT $\boldsymbol{\pi}_{0}$ w/o human H-RDT
@@
H-RDT achieves the highest average success rate of 68.7% in Easy mode and 25.6% in Hard mode, demonstrating significant improvements over existing methods. H-RDT substantially surpasses training from scratch (w/o human) by 8.4% in both Easy and Hard modes, validating the effectiveness of human manipulation pre-training.
@@
Task RDT $\boldsymbol{\pi}_{0}$ w/o human H-RDT
@@
In multi-task settings, H-RDT achieves an average success rate of 87.2%, significantly outperforming RDT (28.8%), $\boldsymbol{\pi}_{0}$ (48.4%), and H-RDT w/o human (67.2%). H-RDT shows a substantial 20.0% absolute improvement over training from scratch, which is notably larger than the improvements observed in single-task scenarios, demonstrating that human manipulation pre-training provides even greater advantages when learning across diverse tasks simultaneously.
@@
Impact of Human Pre-training: The consistent improvements of H-RDT over the baseline without human data across all experimental settings validate our core hypothesis that human manipulation data provides valuable inductive biases. The benefits are most pronounced in few-shot real-world scenarios, where human priors about object affordances and manipulation strategies prove crucial.
@@ Conclusion
Comprehensive evaluations demonstrate consistent improvements over state-of-the-art methods, validating that human manipulation priors provide powerful inductive biases for sample-efficient robotic learning.
@@ A.2 Franka-Panda Detailed Results
Task RDT $\boldsymbol{\pi}_{0}$ w/o human H-RDT
@@ B.4 Real-World Task Definitions
Task ACT DP RDT $\boldsymbol{\pi}_{0}$ w/o human H-RDT